# Synthetic-Powered Predictive Inference

### v1

@@ I.1.2 Results for SPPI with FLUX-generated synthetic data
Figure S11: Results for the ImageNet data using FLUX-generated synthetic images: Coverage rates of OnlyReal, OnlySynth, and SPPI run at level $\alpha=0.02$ (a), $0.05$ (b), and $0.1$ (b), averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Tables S5, S6 and S7 for results across all classes.
Table S5: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials using FLUX-generated synthetic data. The target coverage level is $0.981-\alpha=$ 0.98. The theoretical coverage guarantees for SPPI are in the range $[93.7,100]$. Standard errors are shown in parentheses. Other experimental details follow Figure S11.
@@
Class Only Real Only Synth SPPI Only Real Only Synth SPPI
@@
Table S6: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials using FLUX-generated synthetic data. The target coverage level is $0.951-\alpha=$ 0.95. The theoretical coverage guarantees for SPPI are in the range $[93.7,100]$. Standard errors are shown in parentheses. Other experimental details follow Figure S11.
@@
Class Only Real Only Synth SPPI Only Real Only Synth SPPI
@@
Table S7: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials using FLUX-generated synthetic data. The target coverage level is $0.91-\alpha=$ 0.9. The theoretical coverage guarantees for SPPI are in the range $[81.2,93.7]$. Standard errors are shown in parentheses. Other experimental details follow Figure S11.
@@
Class Only Real Only Synth SPPI Only Real Only Synth SPPI
@@ I.2 Experiments with auxiliary labeled data
Figure S12 presents the results for both marginal and label-conditional guarantees at levels $\alpha=0.05$ and $0.1$, demonstrating trends similar to those observed in Figure 2. The standard conformal prediction, OnlyReal, conservatively controls coverage at the target level $1-\alpha$, but results in larger and noisier prediction sets due to the limited sample size. In contrast, both SPPI-Whole and SPPI-Subset substantially reduce the size and variance of the prediction sets and, as expected, achieve coverage within the theoretical bounds.
Notably, for the “American robin” and “Torch” classes, the SPPI-Subset variant achieves coverage more tightly aligned with the target level $1-\alpha$, outperforming the standard SPPI-Whole method.
We include the results for all real classes in Tables S8, S9 and S10, corresponding to Figures 2, 12(a) and 12(b), respectively.
@@
Figure S12: Results for the ImageNet data: Coverage rates of OnlyReal, SPPI-Whole, and SPPI-Subset run at level $\alpha=0.05$ (a) and $0.1$ (b), averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Tables S9 and S10 for results across all classes.
Table S8: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.981-\alpha=$ 0.98. The theoretical coverage guarantees for both SPPI-Whole and SPPI-Subset are in the range $[93.7,100]$. Other details are as in Figure 2.
@@
Class Only Real SPPI Whole SPPI Subset Only Real SPPI Whole SPPI Subset
@@
Table S9: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.951-\alpha=$ 0.95. The theoretical coverage guarantees for both SPPI-Whole and SPPI-Subset are in the range $[93.7,100]$. Other details are as in Figure S12.
@@
Class Only Real SPPI Whole SPPI Subset Only Real SPPI Whole SPPI Subset
@@
Table S10: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.91-\alpha=$ 0.9. The theoretical coverage guarantee for both SPPI-Whole and SPPI-Subset are in the range $[81.2,93.7]$. Other details are as in Figure S12.
@@
Class Only Real SPPI Whole SPPI Subset Only Real SPPI Whole SPPI Subset
@@
I.2.1 Results for SPPI-Subset with different hyperparameter values
In this section, we present results for the SPPI-Subset procedure across different values of $k$, the number of subsets selected to construct the synthetic calibration set. We compare the performance of SPPI-Subset with SPPI-Whole—which uses all 100 synthetic classes—and the standard conformal prediction, OnlyReal.
Figure S13 presents the performance of all methods for the “American robin" class as a function of $k$, at different values of the level $\alpha$. Notably, for all values of $k$ and $\alpha$, SPPI-Subset achieves coverage within the theoretical bounds.
The two methods, SPPI-Subset and SPPI-Whole, coincide when $k=100$, as both use the full synthetic calibration set. However, for smaller values of $k$, the two methods exhibit significant differences. While SPPI-Whole tends to produce more conservative prediction sets, the SPPI-Subset procedure more tightly achieves the target coverage level across different settings.
For the case $\alpha=0.02$ and $k=5$, both the theoretical lower and upper bounds on coverage are equal to unity, implying that SPPI-Subset yields trivial prediction sets that include all possible classes. This outcome is known a priori and can be avoided by selecting a different hyperparameter for window construction.
@@ I.2.1 Results for SPPI-Subset with different hyperparameter values
Figure S13: Results for the ImageNet data: Coverage rate for OnlyReal,SPPI-Whole, and SPPI-Subset on the American robin class as a function of the number of subsets $k$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
Figure S14 shows the results for the “Beaver” class. For $\alpha=0.02$ and $0.05$, we observe the same trend as in Figure S13: SPPI-Whole yields relatively conservative coverage, while SPPI-Subset with $k<100$ achieves coverage closer to the nominal level $1-\alpha$.
For $\alpha=0.1$, SPPI-Whole—which uses the full synthetic set—already achieves coverage close to the target level $1-\alpha$, suggesting that the empirical $(1-\alpha)$th quantile of the synthetic data closely matches that of the real data. Consequently, in this setting, using only a subset of the synthetic data results in an increase in the variance of the coverage rate.
@@
Figure S14: Results for the ImageNet data: Coverage rate for OnlyReal, SPPI-Whole, and SPPI-Subset on the Beaver class as a function of the number of subsets $k$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
@@
Figure S15 reports the coverage rates and prediction interval lengths for all age groups, evaluated at $\alpha=0.02$, and $0.05$. As in the main paper, we observe that SPPI achieves coverage rates that remain within the theoretical bounds, with lower variance compared to OnlyReal.
@@
Figure S15: MEPS dataset results: coverage and interval length for each age group, obtained by OnlyReal, OnlySynth, and SPPI, at target coverage levels $0.981-\alpha=$ 0.98 (a), and $0.95$ (b). Experiments are repeated over 100 trials. OnlyReal produces trivial (infinite) prediction intervals; thus, its interval length is omitted.
@@
In contrast, the proposed method, SPPI, achieves coverage within the theoretical bounds, closely matching the target $1-\alpha$ level while reducing coverage variance and producing smaller, more informative prediction intervals.
@@
Figure S16: MEPS dataset results: coverage and interval length for the 0–20 age group, obtained by OnlyReal, OnlySynth, and SPPI, at target coverage levels $0.981-\alpha=$ 0.98 (a), $0.95$ (b), and $0.9$ (c). Experiments are repeated over 100 trials. For $\alpha=0.02$ and $0.05$, methods that produce trivial (infinite) prediction intervals are omitted from the interval length panel.
@@
Figure S17: MEPS dataset results: coverage and interval length for the 20–40 age group, obtained by OnlyReal, OnlySynth, and SPPI, at target coverage levels $0.981-\alpha=$ 0.98 (a), $0.95$ (b), and $0.9$ (c). Experiments are repeated over 100 trials. For $\alpha=0.02$ and $0.05$, methods that produce trivial (infinite) prediction intervals are omitted from the interval length panel.

### v2

@@ I.1.2 Results for SPI with FLUX-generated synthetic data
Figure S9: Results for the ImageNet data using FLUX-generated synthetic images: Coverage rates of OnlyReal, OnlySynth, and SPI run at level $\alpha=0.02$ (a), $0.05$ (b), and $0.1$ (b), averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Tables S5, S6 and S7 for results across all classes.
Table S5: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials using FLUX-generated synthetic data. The target coverage level is $0.981-\alpha=$ 0.98. The theoretical coverage guarantees for SPI are in the range $[93.7,100]$. Standard errors are shown in parentheses. Other experimental details follow Figure S9.
@@
Class Only Real Only Synth SPI Only Real Only Synth SPI
@@
Table S6: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials using FLUX-generated synthetic data. The target coverage level is $0.951-\alpha=$ 0.95. The theoretical coverage guarantees for SPI are in the range $[93.7,100]$. Standard errors are shown in parentheses. Other experimental details follow Figure S9.
@@
Class Only Real Only Synth SPI Only Real Only Synth SPI
@@
Table S7: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials using FLUX-generated synthetic data. The target coverage level is $0.91-\alpha=$ 0.9. The theoretical coverage guarantees for SPI are in the range $[81.2,93.7]$. Standard errors are shown in parentheses. Other experimental details follow Figure S9.
@@
Class Only Real Only Synth SPI Only Real Only Synth SPI
@@ I.2 Experiments with auxiliary labeled data
Figure S10 presents the results for both marginal and label-conditional guarantees at levels $\alpha=0.05$ and $0.1$, demonstrating trends similar to those observed in Figure 5. The standard conformal prediction, OnlyReal, conservatively controls coverage at the target level $1-\alpha$, but results in larger and noisier prediction sets due to the limited sample size. In contrast, both SPI-Whole and SPI-Subset substantially reduce the size and variance of the prediction sets and, as expected, achieve coverage within the theoretical bounds.
Notably, for the “American robin” and “Torch” classes, the SPI-Subset variant achieves coverage more tightly aligned with the target level $1-\alpha$, outperforming the standard SPI-Whole method.
We include the results for all real classes in Tables S8, S9 and S10, corresponding to Figures 5, 10(a) and 10(b), respectively.
@@
Figure S10: Results for the ImageNet data: Coverage rates of OnlyReal, SPI-Whole, and SPI-Subset run at level $\alpha=0.05$ (a) and $0.1$ (b), averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Tables S9 and S10 for results across all classes.
Table S8: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.981-\alpha=$ 0.98. The theoretical coverage guarantees for both SPI-Whole and SPI-Subset are in the range $[93.7,100]$. Other details are as in Figure 5.
@@
Class Only Real SPI Whole SPI Subset Only Real SPI Whole SPI Subset
@@
Table S9: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.951-\alpha=$ 0.95. The theoretical coverage guarantees for both SPI-Whole and SPI-Subset are in the range $[93.7,100]$. Other details are as in Figure S10.
@@
Class Only Real SPI Whole SPI Subset Only Real SPI Whole SPI Subset
@@
Table S10: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.91-\alpha=$ 0.9. The theoretical coverage guarantee for both SPI-Whole and SPI-Subset are in the range $[81.2,93.7]$. Other details are as in Figure S10.
@@
Class Only Real SPI Whole SPI Subset Only Real SPI Whole SPI Subset
@@
I.2.1 Results for SPI-Subset with different hyperparameter values
In this section, we present results for the SPI-Subset procedure across different values of $k$, the number of subsets selected to construct the synthetic calibration set. We compare the performance of SPI-Subset with SPI-Whole—which uses all 100 synthetic classes—and the standard conformal prediction, OnlyReal.
Figure S11 presents the performance of all methods for the “American robin" class as a function of $k$, at different values of the level $\alpha$. Notably, for all values of $k$ and $\alpha$, SPI-Subset achieves coverage within the theoretical bounds.
The two methods, SPI-Subset and SPI-Whole, coincide when $k=100$, as both use the full synthetic calibration set. However, for smaller values of $k$, the two methods exhibit significant differences. While SPI-Whole tends to produce more conservative prediction sets, the SPI-Subset procedure more tightly achieves the target coverage level across different settings.
For the case $\alpha=0.02$ and $k=5$, both the theoretical lower and upper bounds on coverage are equal to unity, implying that SPI-Subset yields trivial prediction sets that include all possible classes. This outcome is known a priori and can be avoided by selecting a different hyperparameter for window construction.
@@ I.2.1 Results for SPI-Subset with different hyperparameter values
Figure S11: Results for the ImageNet data: Coverage rate for OnlyReal,SPI-Whole, and SPI-Subset on the American robin class as a function of the number of subsets $k$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
Figure S12 shows the results for the “Beaver” class. For $\alpha=0.02$ and $0.05$, we observe the same trend as in Figure S11: SPI-Whole yields relatively conservative coverage, while SPI-Subset with $k<100$ achieves coverage closer to the nominal level $1-\alpha$.
For $\alpha=0.1$, SPI-Whole—which uses the full synthetic set—already achieves coverage close to the target level $1-\alpha$, suggesting that the empirical $(1-\alpha)$th quantile of the synthetic data closely matches that of the real data. Consequently, in this setting, using only a subset of the synthetic data results in an increase in the variance of the coverage rate.
@@
Figure S12: Results for the ImageNet data: Coverage rate for OnlyReal, SPI-Whole, and SPI-Subset on the Beaver class as a function of the number of subsets $k$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
@@
Figure S13 reports the coverage rates and prediction interval lengths for all age groups, evaluated at $\alpha=0.02$, and $0.05$. As in the main paper, we observe that SPI achieves coverage rates that remain within the theoretical bounds, with lower variance compared to OnlyReal.
@@
Figure S13: MEPS dataset results: coverage and interval length for each age group, obtained by OnlyReal, OnlySynth, and SPI, at target coverage levels $0.981-\alpha=$ 0.98 (a), and $0.95$ (b). Experiments are repeated over 100 trials. OnlyReal produces trivial (infinite) prediction intervals; thus, its interval length is omitted.
@@
In contrast, the proposed method, SPI, achieves coverage within the theoretical bounds, closely matching the target $1-\alpha$ level while reducing coverage variance and producing smaller, more informative prediction intervals.
@@
Figure S14: MEPS dataset results: coverage and interval length for the 0–20 age group, obtained by OnlyReal, OnlySynth, and SPI, at target coverage levels $0.981-\alpha=$ 0.98 (a), $0.95$ (b), and $0.9$ (c). Experiments are repeated over 100 trials. For $\alpha=0.02$ and $0.05$, methods that produce trivial (infinite) prediction intervals are omitted from the interval length panel.
@@
Figure S15: MEPS dataset results: coverage and interval length for the 20–40 age group, obtained by OnlyReal, OnlySynth, and SPI, at target coverage levels $0.981-\alpha=$ 0.98 (a), $0.95$ (b), and $0.9$ (c). Experiments are repeated over 100 trials. For $\alpha=0.02$ and $0.05$, methods that produce trivial (infinite) prediction intervals are omitted from the interval length panel.